Two point-like charges are placed at the corners of a square, a = 34.0 cm. Find the magnitude of the electric field at the bottom left corner. Let q1 = −1.40 µC, and q2 = −3.00 µC. q1 is top right corner and q2 is bottom right corner.
Answer #1

91 E1 12 E2

a34cm0.34m

d- 2 a0.481m

E=kq/r^{2}

-------------------

kg2

E2-9 103 10/(0.34m)

Ег = 233564.01 М/С,

==================

E_{1}=kq_{1}/d^{2}

El = 9 * 10, * 1.4 * 10-6C)(0.481m)

El = 54460 . 35M/C

==================

E, = E2+ E,cos45

E. 233564.0 INCH 54460.35 V/C * cos45

{color{Red} E_{x}=272073.29N/C}

=========================

E, = E, sin45

Ey 54460.35 * sin45

{color{Red} E_{y}=38509.28N/C}

===================

E=sqrt{E_{x}^{2}+E_{y}^{2}}

E=sqrt{(272073.29)^{2}+(38509.28)^{2}}

ANSWER:{color{Red} E=274785.08N/C}

=============================